OpenStudy (anonymous):

-3y + 5 (6y+9) = -1 -5 (4-y)

OpenStudy (anonymous):

-3y+5(6y+9)=-1-5(4-y) -3y+30y+45=-1-20+5y Distribute. 27y+45=-21+5y Combine like terms. -5y -5y Subtract 5y from both sides. 22y+45=-21 -45 -45 Subtract 45 from both sides. 22y=-66 Divide both sides by 22, and the answer is -33.

OpenStudy (anonymous):

that's what i got too, but it is incorrect....

OpenStudy (anonymous):

Does it say it's incorrect in the back of a textbook?

OpenStudy (anonymous):

That's what it said. I've done it so many times and i got so many different answers and put it in, it keep saying incorrect! :/ oh lord.. how much i hate math. sighs*

OpenStudy (anonymous):

Did you copy the original problem down correctly? I sometimes switch positive and negative signs by mistake.

OpenStudy (anonymous):

Yeah, i've tried 33 and -33 it said its both incorrect..